OpenStudy (anonymous):

Simplify: (x+4)*(2x+5)

OpenStudy (anonymous):

help please @iGreen @amistre64

OpenStudy (igreen):

Multiply all the terms in the first parenthesis to all the terms in the 2nd parenthesis.

OpenStudy (igreen):

x * 2x = ? x * 5 = ? 4 * 2x = ? 4 * 5 = ?

OpenStudy (anonymous):

okay so what is X?

OpenStudy (anonymous):

is it 2x squared+13x+20?

OpenStudy (anonymous):

thx but I got the answer tho @iGreen

OpenStudy (igreen):

Yep, you got it.

OpenStudy (anonymous):

yAY!!

OpenStudy (igreen):

\(\Huge\color{#6aaa25}\checkmark\)